OpenStudy (anonymous):

write 480,000 in scientific notation. A rocket travels x0^4 per hour how long would it take for it to reach the moon from the planet.

OpenStudy (unklerhaukus):

\[1=1\times 10^0\]\[10=1\times10^1\]\[100=1\times10^2\]\[1000=1\times10^3\]\[10000=1\times10^4\]\[2100=2.1\times10^3\]

OpenStudy (lgbasallote):

this question doesnt make sense in so many levels =))

OpenStudy (unklerhaukus):

moon rockets dont make sense?

OpenStudy (anonymous):

How would I write 480,000 in scientific notation

OpenStudy (lgbasallote):

well it was asked what the sci notation for 480,000 was...and then it asked for a time given x0^4 which is not defined to be anything...

OpenStudy (anonymous):

I am sorry a typing error I will repeat question

OpenStudy (unklerhaukus):

\[ \text{moon}=\color{yellow}{Yellow^2} \]\[ \text{rocket}=\color{red}{\text{Red}^5} \]

OpenStudy (id21):

\[480,000=48,000\times 10=4800 \times 10^2=480\times 10^3=48 \times 10^4=4.8 \times 10^5\]

OpenStudy (anonymous):

a moon is about 480,0000 miles from the planet it orbits. write 480,000 in scientfic notation. If a rocket travl at 6x10^4 miles per hour, ow long would it take to reach the moon from the planet.

OpenStudy (lgbasallote):

makes more sense

OpenStudy (anonymous):

\[\dfrac {480000} {6\times 10000}\]

OpenStudy (anonymous):

and btw is it 480,000 or 4800,000

OpenStudy (anonymous):

480,000

OpenStudy (anonymous):

assuming it is 480,000 so it is 8 hours